OpenStudy (anonymous):

4x – 9y = - 36 slove for y

OpenStudy (ajprincess):

To solve for y u have to isolate y.i.e u have to have only y on one side of the equation. nw to do that pls follow the steps. 1.Subtract 4x both sides. what do u get when u do so? @karma_258

OpenStudy (anonymous):

subract 4x by 9y or -36?

OpenStudy (ajprincess):

4x-9y-4x=-36-4x

OpenStudy (ajprincess):

4x-9y-4x=-36-4x -9y=-36-4x Nw divide by -9 both sides

OpenStudy (anonymous):

is y= -4/9x+4? just a guess

OpenStudy (ajprincess):